OpenStudy (anonymous):

Solve for x: x = –3 x = 3 x = –3 and x = 3 No Solution

OpenStudy (anonymous):

\[\frac{ 3 }{ 3x } + \frac{ 1 }{ x+4 } = \frac{ 10 }{ 7x }\]

OpenStudy (anonymous):

no solution